OpenStudy (anonymous):

how to solve for lim z->x when 4z2+7z-4x2-7x / z-x

OpenStudy (anonymous):

if you get zero up top, and zero in the bottom (which you do) factor and cancel

OpenStudy (anonymous):

\[4z^2+7z-4x^2-7x\] factor out the \(z-x\) and get \[(z-x)(4x+4z+7)\]

OpenStudy (anonymous):

cancel, replace \(z\) by \(x\)

OpenStudy (anonymous):

this is what I did \[\lim_{z \rightarrow x} 4z ^{2}+7z-4^{2}-7x \z-x\] Then I factor out the z and x \[\lim_{z \rightarrow x} z (4z+7) - x(4x+7) / z-x\] then it becomes: \[\lim_{z \rightarrow x} (z-x) (4z+7)(4x+7( / z-x\] can I cancel the term (z-x) top and bottom now?

OpenStudy (anonymous):

got it, thanks :)

OpenStudy (anonymous):

can I help me solve for one more ??

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

ok wait a second

OpenStudy (anonymous):

i think you factored incorrectly

OpenStudy (anonymous):

same as other, limit problem \[\lim_{z \rightarrow x} 5z+ \sqrt{z} - 5x \sqrt{x} / z-x\]

OpenStudy (anonymous):

as a matter of fact, i am sure you factored incorrectly

OpenStudy (anonymous):

lets go back to the original problem

OpenStudy (anonymous):

which problem? previous or this one? I got the previous one already

OpenStudy (anonymous):

\[\lim_{z \rightarrow x} 5z+\sqrt{z} - 5x - \sqrt{x} / z-x\]

OpenStudy (anonymous):

sorry i got logged off

OpenStudy (anonymous):

this \[z (4z+7) - x(4x+7) \] is incorrect you cannot factor this way

OpenStudy (anonymous):

yes, I got this problem already, I understood how you did this

OpenStudy (anonymous):

because you have no common factors here you have to do something else

OpenStudy (anonymous):

can you help me on the next problem?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[\lim_{z \rightarrow x} 5z + \sqrt{z} - 5x - \sqrt{x} / z-x\]

OpenStudy (anonymous):

\[\lim_{z \rightarrow x}\frac{ 5z+\sqrt{z} - 5x - \sqrt{x} }{z-x}\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

for this one, it is trickier you cannot factor out at \(z-x\) but rather a \((\sqrt{z}-\sqrt{x})\)

OpenStudy (anonymous):

then how would it become?

OpenStudy (anonymous):

it is similar to the first one, assuming you factored correctly (not the one you wrote, but the one i wrote) you can say \[5z-5x+\sqrt{z}-\sqrt{x}=5(\sqrt{z}-\sqrt{x})(\sqrt{z}+\sqrt{x})+(\sqrt{z}-\sqrt{x})\] \[=(\sqrt{z}-\sqrt{x})(5\sqrt{z}+5\sqrt{x}+1)\]

OpenStudy (anonymous):

the denominator also factors as \(z-x=(\sqrt{z}-\sqrt{x})(\sqrt{z}+\sqrt{x})\) so you can cancel the factor of \(\sqrt{z}-\sqrt{x}\)

OpenStudy (anonymous):

after cancellation you get \[\frac{5\sqrt{z}+5\sqrt{x}+1}{\sqrt{z}+\sqrt{x}}\] and then replace \(z\) by \(x\)

OpenStudy (anonymous):

it is very similar to the first one, but instead of factoring the difference of two squares \[z^2-x^2=(z-x)(z+x)\] you factor \(z-x\) as \((\sqrt{z}-\sqrt{x})(\sqrt{z}+\sqrt{x})\)

OpenStudy (anonymous):

I got it, you're so helpful. Can you help me on other problem ?

OpenStudy (anonymous):

sure go ahead and post, but use a new thread

OpenStudy (anonymous):

thanks